# Technology, Institution, and Regional Growth: Evidence from Mineral Mining Industry in Industrializing Japan

### v11

@@ 2.1 Coal Mining Industry
Note: The solid line shows the coal output in giga tons (Gt). The dashed line indicates the coal price in yen per giga tons. The unit price is defined as the total price of coal output divided by the total coal output in each year. Source: Created by the author. Data on coal output are from Ministry of International Trade and Industry (1954).
Figure 1: Coal Production in Japan from 1912 to 1940
@@
In the early 1930s, coal accounted for approximately $55$% of total mineral output in Japan (Mining Bureau, Ministry of Commerce and Industry 1934, p.44). The large share of coal production indicates that coal extraction was an influential economic activity at that time. The higher wage rates of miners have motivated peasants to move into the extraction industry, shifting the local industrial structure from the primary to the production sector. For instance, the average daily wage of female miners working in the pits was one and a half to two times higher than those of the female workers in the silk and spinning industries (Nishinarita 1985, pp. 84–85). Besides the original inhabitants, therefore, mining workers would include a certain proportion of migrants. This may have influenced the structure of both the marriage market and fertility in the local economy surrounding mines. However, quantitative evidence of these changes in the regional economy is still lacking in literature. [^11] Most research on mines in prewar Japan comprises business history studies that investigated the features of the management and industrial organization in coal mining companies (Ishii 1991, pp. 220-233).
@@ 2.2 Mining Technology, Occupational Hazards, and Regulations of Female Miners
By the end of the 19th century, a large part of mines had finished mechanizing the processes involved in hauling, such as drainage pumps and hoisting machines. However, this did not mean that the miner’s workload related to hauling operations was reduced. Instead, such mechanization in the main shafts had increased the workload of miners extracting and transporting coal around the face (Sumiya 1968, pp. 310–312). Moreover, the environment in the pit was generally poor. The air was polluted with dust, oxygen levels were low, and there was a risk of carbon monoxide poisoning. There was also a high fire risk from dynamite blasting and spontaneous combustion of gas and coal. Accidents caused by falling rocks occurred frequently. Available statistics indicate that the average number of accidents per year in the pits between 1917 and 1926, causing injuries and deaths of miners, was 142,595, of which 61,112 were from falling rocks. Accidents caused by explosions, poisoning, and suffocation were fewer than those caused by falling rocks, but they still occurred 311 times per year on average between 1917 and 1926.121212These statistics are from Japan Mining Association (1928, p. table.6).
@@
An important event that changed the harsh labor pattern of female miners was an institutional change that occurred in the early 1930s. Following the International Labor Conference after WWI, the Social Affairs Bureau held a consultation meeting with miners in the 1920s to revise the Miners’ Labor Assistance Regulations of 1916. Initially, mining companies, who relied heavily on female miners for production, opposed the Social Affairs Bureau’s insistence on prohibiting women from working underground and late-night. However, the renewal of mining methods driven by new extraction technologies changed the attitudes of the mining companies in the late 1920s. In fact, the transition from the traditional room-and-pillar mining method to the longwall method was accompanied by the adoption of technologies such as coal cutters and conveyors. Compared to the room-and-pillar method, the coal extraction space of the longwall method is larger, allowing for the introduction of those machines (Nishinarita 1985, p. 91). Specifically, the conveyors could bring extracted coals from pits to gangway further efficiently than the female miners did, adding strong incentive to the mining companies to prohibit the underground works by the female miners (Tanaka and Ogino 1977, p. 79). [^14] Online Appendix A provides finer details in the process of this mechanization.At last, the Revised Miners’ Labor Assistance Regulations of September 1928 prohibited some underground and late-night work by female miners. Since the grace period for enforcement of the regulations lasted until 1933, the number of female miners gradually declined from around 1930.
@@
Panel A of Table 1 shows the evidence that the number of female miners in the pits dropped from 1930 to 1933 when the revised regulation was enacted in September. Despite this, the full enforcement of the revised labor regulations did not eliminate female miners altogether because females were still allowed to work in the pits of the thin layer and were primarily responsible for coal selection (sentan) and other out-of-pit labor. In the late 1930s, more than $15,000$ women still worked in the mines, representing approximately $10$% of all the miners measured. In this light, female miners played an important role in the coal mines throughout the interwar period. The enforcing regulations, however, impacted the market values of female workers in the mining area. Column 1 of Panel B in Table 1 presents evidence that the relative wage rate of female miners dropped after the enforcement. In fact, while there was a decreasing trend in the relative wage rate under the mechanization process in the 1920s, the difference between 1930 and 1933 shows the largest decline in the relative wage rate ($0.66$ v. $0.48$). This reduction is not a macroeconomic trend around the mining region, given that the relative wage rate in the agricultural sector is stable during the same period (Column 2 of Panel B). The rate for in-pit workers dropped from $0.78$ to $0.62$, whereas that for out-pit workers was relatively stable from $0.49$ to $0.45$. This means that regulating in-pit female miners did not improve the relative wage rate of the out-pit female miners. Therefore, although prohibiting female workers in the risky deep layers should mitigate the occupational hazards for female miners, it shall lower the relative wages for females in the mining area.
Table 1: Number of Coal Miners and Relative Wage RateMeasured in the Manufacturing Censuses
@@
Notes: The statistics on the miners in this table are from the Labor Statistics Field Survey Reports that surveyed mines employing 50 or more miners. ‘share of female miners’ in Panel A shows the percentage share of females to the total number of miners. The relative wage rate in Panel B is the female average daily wage divided by the male average daily wage. The overall ratios for the miners between 1927 and 1936 (column 1) are recalculated using the number of miners worked in and outside the pits and their daily wages. Sources: Statistics are from Statistics Bureau of the Cabinet (1926, pp. 5; 32), Statistics Bureau of the Cabinet (1930, pp. 6–7; 26; 30), Statistics Bureau of the Cabinet (1932b, pp. 6–7; 26; 30; 38), Statistics Bureau of the Cabinet (1936, pp. 41; 64; 68; 76), and Statistics Bureau of the Cabinet (1937, pp. 45; 78; 79). The average daily wages for the daily and annual agricultural laborers are obtained from Mataji Umemura, Saburo Yamada, Yujiro Hayami, Nobukiyo Takamatsu, Minoru Kumazaki (1966, p. 107). The description of the survey subject is from Statistics Bureau of the Cabinet (1932a, p. 1) and Statistics Bureau of the Cabinet (1937, p. 1).
@@ 2.3 Mining Technology and Pollutions
First was the soot and smoke caused by coal consumption to run steam-winding machines. Steam-winding engines were introduced early in the mechanization of the hauling process and exhausted the smoke from burning coal in the boiler. However, these steam-powered winders were gradually replaced with electric winders during the interwar period (Japan Mining Association 1930, 1931). Thus, this channel might be unlikely to impact the regional population health. Second, presumably more relevant, is pollution from wastewater generated during mining. In the late 1930s, as coal mining became increasingly mechanized, some mines began introducing water-washing machines during coal sorting. Unfortunately, there is little historical documentation systematically describing pollution caused by wastewater from coal mines. However, coal sludge from coal selection might have contaminated rivers to some extent at that time. An example is Onga river in the Chikuhō coalfield, which was called a “zenzai” (sweet red-bean soup) because the water was contaminated by the liquid waste (Chiba and Yamada 1964). Coal sludge contains several metal toxicities, including mercury (Hg) and cadmium (Cd), which are known to be associated with the incidence of miscarriages and stillbirths (Amadi et al. 2017; Sergeant et al. 2022). Female miners who had worked in the pits before the prohibition of underground works in the early 1930s carried out this coal selection process. Thus, this suggests the possibility that the early-life mortality would have been higher in the coal mining region with greater river accessibility.
@@ 3 Conceptual Framework
The fertility decision in response to the regulation of female miners described in the previous section would affect the population in the mining area. The decline in the relative wage rate of females due to the regulation decreases the opportunity cost of housework, wheras it also decreases their earnings in the labor markets. While the former mechanism may lead to higher fertility through the substitution effect from consumption, the latter may result in lower fertility via the income effect. This section offers a brief overview of a theoretical framework predicting the relevant mechanism given the historical evidence about the coal mines in prewar Japan. To do so, I consider the collective model proposed by Siegel (2017) that illustrates the optimal fertility based on the couple household’s optimization. [^16] For the sake of brevity, I consider a simplified model by abstaining the utility from leisure and home production parameters. This offers a materially similar interpretation of the optimal fertility choice to the original generalized model.
Let us consider that both males and females derive utilities from consumption ($c$) and having children ($b$) under additively separable utility function as $u_{g}(c_{g},b)=\log c_{g}+\tau\log b$, where $j\in\{m,f\}$ indicates a spouse. Raising children requires a home production $y(b)=\pi b$ with $\pi>0$, which is accomplished using labor-saving technologies available in the markets ($s$) and parents’ total home labor input ($\Omega$) as: $y(b)=s^{\gamma}\Omega^{1-\gamma}$. $\gamma$ represents the availability of childcare services in the economy ($10\leq\gamma<$ 1). The parents’ total home labor input is specified in the constant elasticity of the substitution function as: $\Omega=\left(\kappa_{m}h_{m}^{1-\rho}+\kappa_{f}h_{f}^{1-\rho}\right)^{\frac{1}{1-\rho}}$, where $\kappa_{g}$ is the productivity of housework including childcare, $h_{g}\geq 0$ is the home labor input, and $\rho>0$ is the inverse of the elasticity of substitution. The labor income is derived as the product between wage $w_{g}$ and labor input $n_{g}\geq 0$, which is used for their consumption and for purchasing home labor-saving services.
@@
Online Appendix B provides details of the derivation. Under a paternalistic social norm stipulated in the Old Civil Code, the head’s housework productivity, particularly childrearing, shall be negligible ($\kappa_{m}=0$). It follows that condition 1 implies $\frac{w_{f}}{w_{m}}<\frac{1-\gamma}{\gamma}$ for our empirical setting. Since the marketization of childcare ($\gamma$) was considerably low at that time, the optimal fertility is negatively associated with the relative wage ($w_{f}/w_{m}$). [^17] As mechanization progressed, miners were required to accumulate knowledge and experience in operating machinery. Management, therefore, had an incentive to encourage miners to form families and continue employment over a long term (Morimoto 2013). For example, many coal mines were equipped with nurseries and other welfare facilities from the 1920s (Ministry of Commerce and Industry 1926). However, evidence shows that these nurseries required fees and had covered only a small part of their childcares. See Online Appendix B for details. This result suggests the following two predictions:
@@ 4.1 Mine Deposit
Notes: The white circles indicate the location of municipalities with coal mines. The treatment group highlighted in red includes the municipalities within 5 kilometers of a mine. The control group highlighted in pink includes the municipalities between 5 and 30 kilometers from a mine. The excluded municipalities are shown as empty lattices in the figures. Source: Created by the author.
@@
Figure 3: Location of Coal Mines around Chikuhō Coalfield in Kyushū
@@
Notes: 1. The white circles indicate the location of municipalities with coal mines around Chikuhō coalfield (Figure 2a) in Kyushū region. Figure 3a shows the coal mines in Chikuhō, Kasuya, and Miike coalfields in Fukuoka prefecture, Karatsu coalfield in Saga prefecture, and Miike coalfield in Kumamoto prefecture. Amakusa coalfield in Kumamoto prefecture is added in Figure 3b (on the lower left of the figure). Ōita prefecture does not have coal mines but is shown in the figures to explain the border of the sample. 2. Treatment (control) group includes municipalities within 5 (between 5 and 30) km from a mine. The excluded municipalities are shown as empty lattices in the figures. Source: Created by the author.
@@ 4.3 Dependent Variables
To investigate the natural change mechanism behind the local population growth due to mining, I also digitized the Municipal Vital Statistics of 1930 and 1935 (Statistics Bureau of the Cabinet 1932c, 1938), which document the total number of marriages, live births, and deaths in all municipalities in the survey years. I then consider four dependent variables: crude marriage rate, crude birth rate, marital fertility rate, and gender-specific death rate. [^25] I followed the conventional definitions of these variables. The crude marriage and birth rates are the number of marriages and live births per $1,000$ people, respectively. Given the limited data availability on the age-specific population, marital fertility is defined as the number of live births per $1,000$ females. The crude death rate is the number of deaths per $1,000$ people. These are used to test the fertility responses of the miners and the changes in mortality due to mining activities and regulations. To further delve into the occupational hazard channel, I consider the infant mortality rate and fetal death rate. [^26] The infant mortality rate is the number of infant deaths per $1,000$ live births. The number of censored observations in infant mortality is negligible. The fetal death rate is defined as the number of fetal deaths per $1,000$ births. The censored observations in the fetal death rates are less than $11$% in both census years. The estimates from the Tobit estimator (Tobin 1958) are materially similar to those from the OLSE (not reported). This confirms the evidence that the censoring does not matter in my empirical setting. These rates allow for testing the impacts of the occupational hazards for female miners on the mortality selection before birth under the biological framework. I also consider the mortality rate of children aged 1–4 to assess the potential impacts of air pollution. [^27] The child mortality rate is defined as the number of deaths of children aged 1–4 in 1933 (or 1938) per $1,000$ children aged 1–5 in 1930. Note that the number of children aged 1–5 from 1930 is used as it is only documented in the 1930 population census. Given that child deaths, especially at age 5, were rarer than infant deaths at that time, the mismatch in the age bins between numerator and denominator should not be a critical issue. The difference in the survey points does add noise to the estimation of the standard error. However, the coefficient estimate is close enough to zero and thus, should not be a practical issue in this setting (Table 5). To obtain both variables, I digitized the 1933 and 1938 reports documenting municipal-level statistics on births and infant deaths published by the Aiikukai and Social Welfare Bureau.

### v12

@@ 2.1 Coal Mining Industry
Note: The solid line shows the coal output in giga tons (Gt). The dashed line indicates the coal price in yen per giga tons. The unit price is defined as the total price of coal output divided by the total coal output in each year.
Source: Created by the author. Data on coal output are from Ministry of International Trade and Industry (1954).
@@
In the early 1930s, coal accounted for approximately $55$% of total mineral output in Japan (Mining Bureau, Ministry of Commerce and Industry 1934, p.44). The large share of coal production indicates that coal extraction was an influential economic activity at that time. The higher wage rates of miners have motivated peasants to move into the extraction industry, shifting the local industrial structure from the primary to the production sector. For instance, the average daily wage of female miners working in the pits was one and a half to two times higher than those of the female workers in the silk and spinning industries (Nishinarita 1985, pp. 84–85). Besides the original inhabitants, therefore, mining workers would include a certain proportion of migrants. This may have influenced the structure of both the marriage market and fertility in the local economy surrounding mines. However, quantitative evidence of these changes in the regional economy is still lacking in literature. [^11] Most research on mines in prewar Japan comprises business history studies that investigated the features of the management and industrial organization in coal mining companies (Ishii 1991, pp. 220-233). Tanaka (1984) provided a comprehensive review of the works in the Japanese coal mining industry.
@@ 2.2 Mining Technology, Occupational Hazards, and Regulations of Female Miners
By the end of the 19th century, a large part of mines had finished mechanizing the processes involved in hauling, such as drainage pumps and hoisting machines. However, this did not mean that the miner’s workload related to hauling operations was reduced. Instead, such mechanization in the main shafts had increased the workload of miners extracting and transporting coal around the face (Sumiya 1968, pp. 310–312). Moreover, the environment in the pit was generally poor. The air was polluted with dust, oxygen levels were low, and there was a risk of carbon monoxide poisoning. There was also a high fire risk from dynamite blasting and spontaneous combustion of gas and coal. Accidents caused by falling rocks occurred frequently. Available statistics indicate that the average number of accidents per year in the pits between $1917$ and $1926$, causing injuries and deaths of miners, was $142,595$, of which $61,112$ were from falling rocks. Accidents caused by explosions, poisoning, and suffocation were fewer than those caused by falling rocks, but they still occurred $311$ times per year on average between $1917$ and $1926$. [^12] These statistics are from Japan Mining Association (1928, p. table.6).
@@
An important event that changed the harsh labor pattern of female miners was an institutional change that occurred in the early 1930s. Following the International Labor Conference after WWI, the Social Affairs Bureau held a consultation meeting with miners in the 1920s to revise the Miners’ Labor Assistance Regulations of 1916. Initially, mining companies, who relied heavily on female miners for production, opposed the Social Affairs Bureau’s insistence on prohibiting women from working underground and late-night. However, the renewal of mining methods driven by new extraction technologies changed the attitudes of the mining companies in the late 1920s. In fact, the transition from the traditional room-and-pillar mining method to the longwall method was accompanied by the adoption of technologies such as coal cutters and conveyors. Compared to the room-and-pillar method, the coal extraction space of the longwall method is larger, allowing for the introduction of those machines (Nishinarita 1985, p. 91). Specifically, the conveyors could bring extracted coals from pits to gangway further efficiently than the female miners did, adding strong incentive to the mining companies to prohibit the underground works by the female miners (Tanaka and Ogino 1977, p. 79). [^14] Online Appendix A.1 provides finer details in the process of this mechanization. Online Appendix A.2 provides cross-sectional evidence on the replacement of female miners with the use of machines.At last, the Revised Miners’ Labor Assistance Regulations of September 1928 prohibited some underground and late-night work by female miners. Since the grace period for enforcement of the regulations lasted until 1933, the number of female miners gradually declined from around 1930. This change induced by mechanization was commonly observed in Japanese coal mines (Tanaka 1984, p. 392). [^15] A detailed chronology of technology adoptions in each coal mine in Chikuhō coalfield confirms this movement (Compilation Committee of the Chronology of Chikuhō Coal Mining Industry 1973).
@@
Panel A of Table 1 shows the evidence that the number of female miners in the pits dropped from 1930 to 1933 when the revised regulation was enacted in September. Despite this, the full enforcement of the revised labor regulations did not eliminate female miners altogether because females were still allowed to work in the pits of the thin layer and were primarily responsible for coal selection (sentan) and other out-of-pit labor. In the late 1930s, more than $15,000$ women still worked in the mines, representing approximately $10$% of all the miners measured. In this light, female miners played an important role in the coal mines throughout the interwar period. The enforcing regulations, however, impacted the market values of female workers in the mining area. Column 1 of Panel B in Table 1 presents evidence that the relative wage rate of female miners dropped after the enforcement. In fact, while there was a decreasing trend in the relative wage rate under the mechanization process in the 1920s, the difference between 1930 and 1933 shows the largest decline in the relative wage rate ($0.66$ v. $0.48$). This reduction is not a macroeconomic trend around the mining region, given that the relative wage rate in the agricultural sector is stable during the same period (Column 2 of Panel B). The rate for in-pit workers dropped from $0.78$ to $0.62$, whereas that for out-pit workers was relatively stable from $0.49$ to $0.45$. The mean difference before and after the enforcement year is statistically significant only in the in-pit workers (bottom of Panel B). This suggests that regulating in-pit female miners did not improve the relative wage rate of the out-pit female miners. Therefore, although prohibiting female workers in the risky deep layers should mitigate the occupational hazards for female miners, it shall lower the relative wages for females in the mining area.
Table 1: Number of Coal Miners and Relative Wage Rate
Measured in the Manufacturing Censuses
@@
Mean diff. 0.27 0.16 0.05 0.01 0.01
($p$-value) (0.016) (0.046) (0.124) (0.423) (0.769)
Notes: The statistics on the miners in this table are from the Labor Statistics Field Survey Reports that surveyed mines employing 50 or more miners. % share of female miners in Panel A shows the percentage share of females to the total number of miners. The relative wage rate in Panel B is the female average daily wage divided by the male average daily wage. The overall ratios for the miners between 1927 and 1936 (column 1) are recalculated using the number of miners worked in and outside the pits and their daily wages. The mean difference in the wage ratio around 1933 and $t$-statistic $p$-value for the mean difference are listed at the bottom of Panel B.
Sources: Statistics are from Statistics Bureau of the Cabinet (1926, pp. 5; 32), Statistics Bureau of the Cabinet (1930, pp. 6–7; 26; 30), Statistics Bureau of the Cabinet (1932b, pp. 6–7; 26; 30; 38), Statistics Bureau of the Cabinet (1936, pp. 41; 64; 68; 76), and Statistics Bureau of the Cabinet (1937, pp. 45; 78; 79). The average daily wages for the daily and annual agricultural laborers are obtained from Umemura et al. (1966, p. 107). The description of the survey subject is from Statistics Bureau of the Cabinet (1932a, p. 1) and Statistics Bureau of the Cabinet (1937, p. 1).
@@ 2.3 Mining Technology and Pollutions
First was the soot and smoke caused by coal consumption to run steam-winding machines. Steam-winding engines were introduced early in the mechanization of the hauling process and exhausted the smoke from burning coal in the boiler. However, these steam-powered winders were gradually replaced with electric winders during the interwar period (Japan Mining Association 1930, 1931). Thus, this channel might be unlikely to impact the regional population health. Second, presumably more relevant, is pollution from wastewater generated during mining. As coal mining became increasingly mechanized, some mines began introducing water-washing machines during coal sorting. Unfortunately, there is little historical documentation systematically describing pollution caused by wastewater from coal mines. However, coal sludge from coal selection had contaminated rivers to some extent at that time. An example is Onga river in the Chikuhō coalfield, which was called a “zenzai” (sweet red-bean soup) because the water was contaminated by the liquid waste (Chiba and Yamada 1964). Coal sludge contains several metal toxicities, including mercury (Hg) and cadmium (Cd), which are known to be associated with the incidence of miscarriages and stillbirths (Amadi et al. 2017; Sergeant et al. 2022). This suggests the possibility that early-life mortality would have been higher in the coal mining region with greater river accessibility. From the 1920s, however, some mines installed settling ponds to purify wastewater and circulate it to coal cleaning (Advisory Council on Mining 1932, pp. 313–316). This might work to mitigate the pollution risks around the mines due to the wastewater.
@@ 3 Conceptual Framework
The fertility decision in response to the regulation of female miners described in the previous section would affect the population in the mining area. The decline in the relative wage rate of females due to the regulation decreases the opportunity cost of housework, wheras it also decreases their earnings in the labor markets. While the former mechanism may lead to higher fertility through the substitution effect from consumption, the latter may result in lower fertility via the income effect. This section offers a brief overview of a theoretical framework predicting the relevant mechanism given the historical evidence about the coal mines in prewar Japan. To do so, I consider the collective model proposed by Siegel (2017) that illustrates the optimal fertility based on the couple household’s utility maximization. [^17] For the sake of brevity, I consider a simplified model by abstaining the utility from leisure and home production parameters. This offers a materially similar interpretation of the optimal fertility choice to the original generalized model. See Browning et al. (2014) for a comprehensive review on the cooperative and non-cooperative models of the couples.
Let us consider that both males and females derive utilities from consumption ($c$) and having children ($b$) under additively separable utility function as $u_{g}(c_{g},b)=\log c_{g}+\tau\log b$, where $g\in\{m,f\}$ indicates a spouse. Raising children requires a home production $y(b)=\pi b$ with $\pi>0$, which is accomplished using labor-saving technologies available in the markets ($s$) and parents’ total home labor input ($\Omega$) as: $y(b)=s^{\gamma}\Omega^{1-\gamma}$. $\gamma$ represents the availability of childcare services in the economy ($10\leq\gamma<$ 1). The parents’ total home labor input is specified in the constant elasticity of the substitution function as: $\Omega=\left(\kappa_{m}h_{m}^{1-\rho}+\kappa_{f}h_{f}^{1-\rho}\right)^{\frac{1}{1-\rho}}$, where $\kappa_{g}$ is the productivity of housework including childcare, $h_{g}\geq 0$ is the home labor input, and $\rho>0$ is the inverse of the elasticity of substitution. The labor income is derived as the product between wage $w_{g}$ and labor input $n_{g}\geq 0$, which is used for their consumption and for purchasing home labor-saving services.
@@
Online Appendix B provides details of the derivation. Under a paternalistic social norm stipulated in the Old Civil Code, the head’s housework productivity, particularly childrearing, shall be negligible ($\kappa_{m}=0$). It follows that condition 1 implies $\frac{w_{f}}{w_{m}}<\frac{1-\gamma}{\gamma}$ for my empirical setting. Since the marketization of childcare ($\gamma$) was considerably low at that time, the optimal fertility is negatively associated with the relative wage ($w_{f}/w_{m}$). [^18] As mechanization progressed, miners were required to accumulate knowledge and experience in operating machinery. Management, therefore, had an incentive to encourage miners to form families and continue employment over a long term (Ogino 1993). For example, many coal mines were equipped with nurseries and other welfare facilities from the 1920s (Ministry of Commerce and Industry 1926). However, evidence shows that these nurseries required fees and had covered only a small part of their childcares. See Online Appendix B for details. This result suggests the following two predictions:
@@ 4.1 Mine Deposit
Notes: The white circles indicate the location of municipalities with coal mines. The treatment group highlighted in red includes the municipalities within 5 kilometers of a mine. The control group highlighted in pink includes the municipalities between 5 and 30 kilometers from a mine. The excluded municipalities are shown as empty lattices in the figures.
Source: Created by the author.
@@
Notes: 1. The white circles indicate the location of municipalities with coal mines around Chikuhō coalfield (Figure 2a) in Kyushū region. Figure 3a shows the coal mines in Chikuhō, Kasuya, and Miike coalfields in Fukuoka prefecture, Karatsu coalfield in Saga prefecture, and Miike coalfield in Kumamoto prefecture. Amakusa coalfield in Kumamoto prefecture is added in Figure 3b (on the lower left of the figure). Ōita prefecture does not have coal mines but is shown in the figures to explain the border of the sample.
2. Treatment (control) group includes municipalities within 5 (between 5 and 30) km from a mine. The excluded municipalities are shown as empty lattices in the figures.
Source: Created by the author.
@@ 4.3 Dependent Variables
To investigate the natural change mechanism behind the local population growth due to mining, I also digitized the Municipal Vital Statistics of 1930 and 1935 (Statistics Bureau of the Cabinet 1932c, 1938), which document the total number of marriages, live births, and deaths in all municipalities in the survey years. I then consider four dependent variables: crude marriage rate, crude birth rate, marital fertility rate, and gender-specific death rate. [^26] I followed the conventional definitions of these variables. The crude marriage and birth rates are the number of marriages and live births per $1,000$ people, respectively. Marital fertility is defined as the number of live births per $1,000$ married females. Although the number of married females in each age is unavailable, the number of married females is similar to the number of households in each municipality, meaning that the marital fertility definition herein is useful in representing the couple’s fertility decision. In other words, the results from this marital fertility definition shall not be sensitive to the existence of the elder married women in the households. Although marital fertility is considered to capture the couple’s fertility decision better, the results can be materially similar between crude and marital fertilities because out-of-wedlock births have been historically rare in Japan. The crude death rate is the number of deaths per $1,000$ people. These are used to test the fertility responses of the miners and the changes in mortality due to mining activities and regulations. To further delve into the occupational hazard channel, I consider the infant mortality rate and fetal death rate. [^27] The infant mortality rate is the number of infant deaths per $1,000$ live births. The number of censored observations in infant mortality is negligible. The fetal death rate is defined as the number of fetal deaths per $1,000$ births. The censored observations in the fetal death rates are less than $11$% in both census years. The estimates from the Tobit estimator (Tobin 1958) are materially similar to those from the OLSE (not reported). This confirms the evidence that censoring does not matter in my empirical setting. These rates allow for testing the impacts of the occupational hazards for female miners on the mortality selection before birth under the biological framework. I also consider the mortality rate of children aged 1–4 to assess the potential impacts of air pollution. [^28] The child mortality rate is defined as the number of deaths of children aged 1–4 in 1933 (or 1938) per $1,000$ children aged 1–5 in 1930. Note that the number of children aged 1–5 from 1930 is used as it is only documented in the 1930 population census. Given that child deaths, especially at age 5, were rarer than infant deaths at that time, the mismatch in the age bins between numerator and denominator should not be a critical issue. The difference in the survey points does add noise to the estimation of the standard error. However, the coefficient estimate is close enough to zero and thus, should not be a practical issue in this setting (Table 5). To obtain both variables, I digitized the 1933 and 1938 reports documenting municipal-level statistics on births and infant deaths published by the Aiikukai and Social Welfare Bureau.